OpenStudy (anonymous):

Find the solution to the following system of equations. x + 3y = 12 x + y = 2

OpenStudy (tkhunny):

Have you considered subtracting the second from the first?

OpenStudy (wolf1728):

Following tkhunny's suggestion: x + 3y = 12 -x - y = -2 Now add the 2 equations.

OpenStudy (wolf1728):

Okay, I'll add the equations and get: 2y = 10 So y equals?

OpenStudy (anonymous):

5?

OpenStudy (tkhunny):

Just for the record, I VERY MUCH recommend multiplication by -1 followed by addition. MUCH better than subtraction.